most air travellers now use e-tickets. Electronic ticketing allows passengers to not worry about a paper ticket, and it costs the airline companies less to handle than paper ticketing. However, in recent times, the airlines have received complaints from passengers regarding their e-tickets, particularly when connecting flights and a change of airlines were involved. To investigate the problem, an independent agency contacted a random sample of 20 airports and collected information on the number of complaints the airport had with e-tickets for the month of March. The information is reported below: 10 14 11 11 10 18 10 16 18 11 18 11 16 18 15 10 13 11 14 10 At the 0.02 significance level, can the agency conclude that the mean number of complaints per airport is less than 15 per month?

a. What assumption is necessary before conducting a test of hypothesis?

Conduct a test of hypothesis and interpret the results. i need this part answer H0 : μ ≥ 15; H1 : μ < 15; H0 if t <..... . (Round the final answer to 3 decimal places.) The value of the test statistic is -(Negative answer should be indicated by a minus sign. Round the final answer to 2 decimal places.)

i just need t value on 0.02
Answer #1

Solution:

Given that sample n =20

Mean x̄:   13.25

Standard Deviation, s: 3.143

H0 : μ ≥ 15

Ha : μ < 15;

crit value for one-tail test, alpha=5%,

df = 20: t =invT(0.95,20)= 1.7247

test statistic: t(13.25) = (13.25-15)/(3.143/sqrt(21)) = -2.55

Conclusion: since the test stat is not in the reject interval, fail
to reject Ho.